OpenStudy (gabylovesyou):

sqrt(-25) / (5-2i)+(1-3i)

OpenStudy (gabylovesyou):

\[\frac{ \sqrt{-25} }{ (5-2i) + (1-3i) }\]

OpenStudy (gabylovesyou):

@jhonyy9 @ikram002p

OpenStudy (gabylovesyou):

@freckles

OpenStudy (gabylovesyou):

is it \[\frac{ -25 +30i }{ 61 }\]

jhonyy9 (jhonyy9):

sqrt(-25) −−−−-------- = (5−2i)+(1−3i) how you have got it ?

jhonyy9 (jhonyy9):

sqrt(-25) = 5i yes ?

OpenStudy (gabylovesyou):

cuz i first got \[- \frac{ 5(5-6i) }{ 61 }\]

OpenStudy (gabylovesyou):

idk how to simplify that..

jhonyy9 (jhonyy9):

let 5-2i = a and 1-3i = b so than in denominator there are a+b but you know what mean conjugate of something ?

OpenStudy (gabylovesyou):

yes.. to turn the symbol around.. so instead of - its +

jhonyy9 (jhonyy9):

hope you know formula a^2 -b^2 = (a-b)(a+b)

OpenStudy (gabylovesyou):

i do

jhonyy9 (jhonyy9):

than multiplie down and up by a-b - what will get ?

OpenStudy (gabylovesyou):

what do you mean ?

jhonyy9 (jhonyy9):

check what i have wrote above

OpenStudy (gabylovesyou):

im confused.

OpenStudy (gabylovesyou):

i dont understand what i have to do

jhonyy9 (jhonyy9):

(5−2i)+(1−3i) - this is the denominator a + b than multiplie down and up by a-b so you will get a^2 -b^2 what mean a^2 =(5-2i)^2 and b^2 = (1-3i)^2 yes ?

OpenStudy (gabylovesyou):

yes i understand .

jhonyy9 (jhonyy9):

distribute parentheses -- see what will get - remain in denominator - what you need to make - try eliminate the terms with i from denominator

OpenStudy (phi):

it looks a bit painful. I would first add the bottom: reals to reals and imaginary to imaginary to get 1 complex number can you do that ?

OpenStudy (gabylovesyou):

you mean like 6 - 5i ?

OpenStudy (gabylovesyou):

@phi

jhonyy9 (jhonyy9):

yes - so will be more easy

jhonyy9 (jhonyy9):

so than with 6-5i multiplie numerator and denominator by conjugate of 6-5i do you know it ?

OpenStudy (gabylovesyou):

conjugate is 6 +5i but what do i multiply it with ?

OpenStudy (phi):

yes

jhonyy9 (jhonyy9):

(6-5i)(6+5i) = ?

OpenStudy (gabylovesyou):

61?

jhonyy9 (jhonyy9):

yes

OpenStudy (phi):

you multiply top and bottom by the conjugate 6+5i the bottom is 36+25 = 61 (looks like that is what you got before) the top is 5i(6+5i) = -25 +30i

jhonyy9 (jhonyy9):

so this will get for denominator

OpenStudy (phi):

it looks like you did it correctly up top

OpenStudy (gabylovesyou):

soo the answer is correct.. its \[\frac{ -25 + 30i }{ 61 }\]

OpenStudy (gabylovesyou):

ok thank you

OpenStudy (phi):

sometimes, people write it as -25/61 + 30/61 i so that it is in real/imag format

jhonyy9 (jhonyy9):

was my pleasure -